OpenStudy (anonymous):

I need help please MEDAL FAN Solve each equation. Write Justification for each step. 5x-3 = 4(x+2)

OpenStudy (anonymous):

5x-3 = 4(x+2) 4*x , 4*2 - Distributive Property 5x-3 = 4x+8 +3 +3 - Addition Property 5x = 4x+11 -4x -4x - Subtraction Property 1x = 11 -- -- - Division Property 1 1 x = 11 <-- Answer

Directrix (directrix):

I agree with your work. Some people may argue that the reason for the last step is the multiplicative property of one.

Directrix (directrix):

@Pepper5

OpenStudy (anonymous):

Thank you both so much!!! :)

Directrix (directrix):

You are welcome.